OpenStudy (campayne):

Classify the system and determine the number of solutions. (Check all that apply. ) {7x - y = -11 3y = 21x + 33 infinitely many solutions independent consistent no solution inconsistent dependent

satellite73 (satellite73):

now it makes more sense

satellite73 (satellite73):

solve both for y, you will see it is the same line

OpenStudy (sushi121212):

It would be easier if you wrote it so they both say Ax+Bx=C: 7x - y = -11 -21x + 3y = 33 Then you can solve for x or y.

OpenStudy (sushi121212):

If the number you get equals each other, then there is an infinite number of equations (like 3=3), and if you get 2 number that do not equal each other (like 5=8), then it will be no solution. If you get y=___ or x=_____ then there is one solution (like x=6, y=3). Consistent is where the numbers are the same (infinite # of solutions) Inconsistent is where there is no solutions. Dependent is where there is the same # on both sides (infinite # of solutions) and independent is when you get x=_____ or y=_____.

OpenStudy (campayne):

okay i think i get it